Suppose that X is a random variable with mean = 10 and standard deviation = 3. Suppose that Y is a random variable with mean = 20 and standard deviation = 4. Suppose that X and Y are independent. Find the standard deviation of X+Y.

A.

5

B.

7

C.

3.5

D.

25
Answer #1

The variance of X+Y is given by,

Var(X+Y) = Var(X) + Var(Y) +2Cov(X,Y)

Since , X and Y are independent.

Therefore , Cov(X,Y) = 0

Therefore ,

Var(X+Y) = Var(X) + Var(Y) +2Cov(X,Y) = 32 + 42 + 2*0 = 25

The standard deviation of X+Y is   =